# Efficient Data Selection for Training Genomic Perturbation Models

### v5

@@ Introduction
Similar to  huang2024sequential , we start with a random test set that contains roughly 10% of the available perturbations. The test set stays constant throughout the cycles to ensure there is no interaction between the strategies and the evaluation, which could lead to inconsistencies, e.g., adding a test perturbation that is hard to predict in the train set would diminish the error due to easier evaluation, not due to the strategy’s success. We train the model 5 times sequentially, each time adding to the train set 5% of the available perturbations and keeping 1% for validation. At the end of the process, 30% is used for training and validation, and 10% for testing, leaving 60% of the dataset unutilized. This is imperative to perform an effective comparison between the selected train sets. To be specific, if we increase the train set size, the methods converge unavoidably to very similar final gene selections because the available choices are limited. Hence we choose to leave enough space between choices to highlight the differences between the strategies.
We use a 10-fold cross validation to define the test sets, and perform the whole experiment for 3 different random seed initializations (including model initialization and data splits).We evaluate stability by quantifying how much do the chosen gene perturbations diverge between different seed initialization for each method. To this end, we plot the number of unique genes covered by each method throughout 3 different initializations. Moreover, we quantify the consistency of each method by measuring the overlap between the resulting gene sets using the average pairwise Jaccard similarity defined in the appendix. The evaluation metric for performance is MSE and is computed on the top 20 most differentially expressed genes, i.e., the genes that exhibit the biggest expression differences in the experiment, akin to the literature roohani2024predicting . This is a common practice since the vast majority of the genes have zero expression, meaning the results would be skewed and the differences minuscule had we computed them globally. We use the default GEARS parameters and implementation [^1] https://github.com/snap-stanford/GEARS/tree/master.
@@ Results
Fig. 3 illustrates the number of unique genes selected to train GEARS for 3 different random initializations and the average Jaccard similarity through all cycles is in Tab. 1 (the full tables can be found in the Appendix). The random seed affects the model and the k-folds of the data. The budget represents the number of genes added per cycle and acts as a baseline. The set of genes retrieved by the data filtering methodologies differs from the budget solely due to different k-fold splits. In contrast, active learning methodologies exhibit near-random variability across different runs. This instability implies that replicating the model would require spending multiple times the original ’gene’ budget, undermining reproducibility. Moreover, the resulting gene sets are tightly coupled to model-specific biases, limiting their reusability for downstream analyses or integration with other studies, and indicating that the selection process is driven more by model bias rather than by a signal in the data.
Data Random
ACS
FW
Typi
Clust
Max
Spec
Graph
Reach
Adamson 0.15 0.15 0.15 0.75 0.75
Replogle 0.10 0.13 0.10 0.78 0.78
Table 1: Average Jaccard similarity on gene selections.
Test Error
As can be seen in Tab. 2, the final model perofrmances of GraphReach are close to TypiClust. Considering the overlapping confidence intervals and similar performance in both datasets, we can deduce that GraphReach performs competitively to the state-of-the-art. In contrast, MaxSpec’s performance in Adamson is inferior. The potential cause is that the Adamson gene sets are limited to 25 nodes, meaning that the spectral norm of the subgraph formed from the respective receptive field might not suffice to act as a proxy for generalization, in contrast to Replogle, where MaxSpec’s performance is strong. Similarly, ACS-FW also performs better in Replogle, possibly due to the increased number of training genes, which makes the model more accurate, as indicated by the MSE. Despite their success in Replogle, both ACS-FW and MaxSpec fall short with the lack of supervision in Adamson, rendering them inferior overall to GraphReach and TypiClust. A final observation is that random selection can be quite effective, albeit with great variance, a fact well known in active learning under budget simeoni2021rethinking ; mittal2019parting ; zhu2019addressing and agreeing with the recent literature on the problem at hand huang2024cost .
Tradeoff
The overall tradeoff between efficiency, test error, and stability is visualized in Fig. 4. We observe that, overall, GraphReach strikes the best balance between low MSE, training efficiency, and stability. We thus deem it a valid alternative to random or active-learning-based training for GEARS, especially given its simple deployment and better overall performance compared to random selection.
Method MSE ($10^{3}$) ↓ AVG
Adamson Replogle
Random $184\pm 13.63$ $159\pm 0.80$ 171.5
ACS-FW $185\pm 9.42$ $154\pm 1.15$ 169.5
TypiClust $180\pm 1.80$ $155\pm 1.21$ 167.5
MaxSpec $192\pm 7.11$ $155\pm 1.47$ 173.5
GraphReach $180\pm 1.93$ $156\pm 1.09$ 168
Table 2: MSE, best is bold and second best is underlined.
@@ Conclusion
Genomic experiments cannot exhaust the available intervention options because they are too costly and time-consuming, thus, GNNs are utilized to predict the outcomes of these interventions (i.e., gene perturbations). However, the training set for a gene expression model consists of genomic experiments themselves, calling for more efficient training strategies. In this work, we proposed graph-based data filtering methods that build the train set without relying on the model’s output. This provides significant real-world speedup (from five months to one) as well as more stable selections. We developed two methods, GraphReach and MaxSpec, which build the train set by selecting the perturbations that are expected to improve the GEARS’ generalization. In the experiments, we observed that GraphReach exhibits competitive test error overall, while being significantly faster and more stable compared to the other methods. In the future, we plan to examine hybrid techniques that utilize data filtering methods to jump-start the training up to a point and blend in active learning for the rest of the runs. Moreover, we plan to experiment with multi-gene perturbations and potentially benchmark with graph active learning methods zhang2021rim ; zhang2021grain .
@@
Appendix A Proof of Submodularity
@@
Appendix B Stability with Jaccard Similarity
The stability is defined as the average pairwise Jaccard similarity between the gene sets that each method chooses between runs with different random seed.
$\bar{J}=\frac{1}{\binom{n}{2}}\sum_{1\leq i<j\leq n}J(S_{i},S_{j})=\frac{2}{n(n-1)}\sum_{1\leq i<j\leq n}\frac{|S_{i}\cap S_{j}|}{|S_{i}\cup S_{j}|}$ (19)
The full results can be seen in Table 3.
Data Cycle 0 1 2 3 4
Adamson
Random 0.13 0.12 0.14 0.16 0.20
ACS-FW 0.12 0.11 0.14 0.18 0.22
TypiClust 0.12 0.12 0.15 0.18 0.22
MaxSpec 0.76 0.74 0.75 0.75 0.76
GraphReach 0.78 0.76 0.74 0.75 0.76
Replogle
Random 0.04 0.07 0.10 0.14 0.17
ACS-FW 0.04 0.08 0.13 0.18 0.24
TypiClust 0.04 0.07 0.10 0.14 0.18
MaxSpec 0.77 0.78 0.79 0.79 0.79
GraphReach 0.78 0.78 0.79 0.79 0.79
Table 3: Jaccard similarity on gene selections per cycle.

### v6

@@ Introduction
For GraphReach and Baseline, the train and validation sets are defined before the acquisition of the single-cell data and they require no interaction with the wet-lab. Thus, they are expected to take around 30 days, which corresponds to the time the PerturbSeq experiment takes, since running GraphReach at the beginning and training GEARS at the end takes negligible time. For all methods, 25% of the gene perturbations are used for training and 5% for validation, 10% are kept for testing, leaving 60% of the dataset unutilized. This is imperative to perform an effective comparison between the selected train sets. If we increase the train set size, the methods converge unavoidably to very similar final gene selections because the available choices are limited. Hence, we leave sufficient space between choices to highlight the differences between the strategies. We use a 10-fold cross validation to adjust the 10% test set and take the average performance. The whole experiment is repeated for 3 different random seed initializations (including model initialization and data splits).
For the active learning methods, in accordance to  (huang2024sequential, ), we keep the initial 10% test set constant throughout the cycles to ensure there is no interaction between the strategies and the evaluation. We train the model a number of times sequentially, each time adding to the train and validation set a percentage of the available perturbations until they reach 25% and 5%, respectively. In order to quantify the relationship between the training cycles and final model accuracy, active learning methods are run for 3 and 5 cycles, the latter being the main choice in the literature. The models with 5 cycles receive 5% of the perturbations in each cycle for training and 1% for validation, while for 3 cycles the numbers are 8.3% and 1.7% respectively. Models trained on 5 cycles are expected to take 150 days, and 3 cycles correspond to 90 days. The methods are evaluated based on the performance of GEARS trained on the final cycle. It should be noted that no active learning method can go less than 60 days, as 1 training cycle (30 days) means no input from the model.
Following the literature, the evaluation metrics for performance are Pearson correlation and MSE, which is computed on the top 20 most differentially expressed genes, i.e., the genes that exhibit the biggest expression differences in the experiment. This is a common practice since the vast majority of the genes have zero expression, meaning the results would be skewed and the differences minuscule had we computed them for more genes (roohani2024predicting, ). We evaluate stability by quantifying how much the chosen gene perturbations diverge between different seed initializations for each method. To this end, we quantify the consistency of each method by measuring the overlap between the resulting gene sets using the average pairwise Jaccard similarity defined in the appendix. To evaluate robustness to noise, we performed the same experiments where GraphReach runs on corrupted versions of the knowledge graph. Specifically, we delete 5% and 10% of the edges and add the same amount of fake ones, to represent a common setting of errors in the graph. In all experiments, we use the default GEARS parameters and implementation [^1] https://github.com/snap-stanford/GEARS/tree/master.
@@ Results
Accuracy
The overall tradeoff between accuracy and efficiency is visualized in Fig. 3. For Pearson correlation, we observe that GraphReach achieves the strongest performance on average, as it performs on par or better than the benchmarks in three of the four datasets, despite taking a fraction of the time. We see that it outperforms the Baseline, the only benchmark with similar efficiency, in 7 out of 8 evaluations, and that the latter is significantly more uncertain in terms of confidence intervals.
On average, the active learning methods perform on par with each other, with TypiClust being better in Adamson and Norman and ACS-FW being better in the rest. Another observation for ACS-FW is that it remains constant for 3 and 5 cycles for Adamson, 3 cycles are better than 5 in Norman while the results are split in K562 and RPE1. We believe that this happens because training on 3-cycles increases the amount of samples the model sees per cycle and this allows GEARS to provide better uncertainty estimates which are integral for ACS-FW. In contrast, TypiClust tends to deteriorate as we constrain the days with only one exception in RPE1 for Pearson correlation. That said, ACS-FW has considerably larger confidence intervals (competing with the Baseline) throughout all experiments, possibly due to the sensitivity to GEARS’ uncertainty estimation. We analyze further this instability in the coming sections. We thus deem TypiClust the second best method, as it is considerably more stable and achieves competitive performance.
Note that, as mentioned above, active learning requires by default a number of cycles to run and hence it can not achieve similar efficiency to GraphReach and we can not compare them head-to-head. However, if we constrain the experiments to up to 90 days, GraphReach performs better than active learning in 5 out of 8 evaluations. Given that on average active learning deteriorates as we constrain the days, we can deduce that GraphReach showcases the best tradeoff between accuracy and efficiency.
@@
The average Jaccard similarity between the perturbation sets selected through all cycles can be found in Tab. 2. The random seed affects the model and the k-folds of the data, hence the set of genes retrieved by GraphReach differ solely due to different splits. In contrast, active learning methodologies exhibit variability across different runs that approximate the Baseline, which is a random selection. This instability implies that the resulting gene sets are tightly coupled to model-specific biases, limiting their reusability for downstream analyses or integration with other studies, and indicating that the selection process is driven more by model bias rather than by a signal in the data.
Table 2: Average Jaccard similarity on gene selections through different random seed initializations.
Data Base line ACS FW Typi Clust Graph Reach
Adamson 0.15 0.15 0.15 0.75
K562 0.10 0.13 0.10 0.78
RPE1 0.10 0.13 0.10 0.76
Norman 0.11 0.13 0.15 0.95
Table 3: Performance under different noise levels in the knowledge graph, as represented by random removal and addition of edges. Results are rounded up to the nearest integer.
Dataset Noise (%) Pearson (↑) MSE (↓)
Adamson 0 $952\pm 1$ $180\pm 2$
5 $951\pm 1$ $181\pm 6$
10 $947\pm 1$ $196\pm 8$
K562 0 $919\pm 1$ $156\pm 1$
5 $919\pm 2$ $154\pm 1$
10 $920\pm 2$ $153\pm 2$
RPE1 0 $706\pm 3$ $128\pm 1$
5 $701\pm 4$ $129\pm 1$
10 $704\pm 6$ $126\pm 2$
Norman 0 $895\pm 5$ $168\pm 10$
5 $893\pm 13$ $170\pm 18$
10 $894\pm 13$ $163\pm 13$
Robustness
The results from the robustness experiments are in Tab. 3. We observe that while GraphReach underperforms in noisy settings in the Adamson dataset, it actually stays close to or even surpasses the original in the rest. We hypothesize that this happens because random edges may have a relatively high probability of connecting previously distant parts of the graph and may therefore further increase the coverage that GraphReach is able to achieve. So, we observe satisfactory empirical robustness of GraphReach. Similar effects of random perturbations on approximation quality have been observed in studies of vertex cover and related graph problems shi2018runtime .
@@ Conclusion
Genomic experiments cannot exhaust the available intervention options, thus, gene perturbation models are used to predict the outcomes of these interventions. However, their training sets consist of costly genomic experiments themselves, calling for efficient training strategies. In this work, we focused on graph neural networks and propose GraphReach, a subset-selection method that chooses perturbations such that the model is trained more efficiently. In the experiments, we observed that compared to existing active learning solutions, GraphReach exhibits significant real-world speedup (from five months to one) as well as more stable selections without sacrificing accuracy in general and while being robust to errors in the knowledge graph. We deem GraphReach a valid alternative to random or active-learning-based training given its simple deployment and better overall performance. In the future, we plan to examine hybrid techniques that utilize subset selection methods to jump-start the training up to a point and blend in active learning for the rest of the runs. Moreover, we plan to research the use of subset-selection for non-graph-based methods, covering both foundation models (theodoris2023transfer, ) and regression (ahlmann2025deep, ).
@@
Appendix A Proof of Proposition 1
We begin by considering the gradient of the loss $\mathcal{L}$ with respect to the identity embeddings $\mathbf{W}_{0}$ as a function of $\frac{\partial\mathcal{L}}{\partial\mathbf{H}}$:
Thus, the full gradient becomes:
which shows that the gradient of $\mathbf{W}_{0}[i]$ depends only on the gradients of its neighbors in $\hat{\mathbf{A}}$. However, $\frac{\partial\mathcal{L}}{\partial\mathbf{H}[j]}=0$ if $j\notin S$.
Appendix B Proof of Proposition 2